# Rational torsion on hyperelliptic jacobian varieties

### v1

@@ Abstract.
Fixing an integer $g\geq 2$, it was conjectured by Flynn that there exists a constant $\kappa$, independent of $g$, such that for every $m\leq\kappa g$, there exists a hyperelliptic curve of genus $g$ over ${\mathbb{Q}}$ with a rational $m$-torsion point on its Jacobian. Leprévost proved this conjecture with $\kappa=3$. In this work we prove that given an integer $N$ in the interval $[3g,4g+1]$, $g\geq 3$, satisfying certain partition conditions, there exist infinite families of hyperelliptic Jacobian varieties with a rational torsion point of order $N$. In particular, we establish the existence of such varieties for $N=4g+1$ when $g$ is odd and for $N=4g-1$ when $g$ is even. A few explicit applications of this result produce the first known infinite examples of torsion $13$ when $g=3$, torsion $15$ when $g=4$, and torsion $17,18,21$ when $g=5$. In fact, we show that infinitely many of the latter abelian varieties are absolutely simple.
@@ 1. Introduction
A computational approach toward studying torsion subgroups of abelian varieties of dimension $d\geq 2$ is to find integers that are realized as orders of torsion points on these varieties. For example, a family of genus $2$ curves whose Jacobian varieties possess torsion points of order $\ell$ over ${\mathbb{Q}}$ was given in [7] when $\ell=11$, and in [13, 14, 16], when $\ell=13$, $15$, $17$, $19$ or $21$. In [9], several explicit families of curves of genus $2$ and $3$ whose Jacobians possess large rational torsion subgroups were displayed.
Researchers questioned the possibility of constructing families of abelian varieties in which the size of the rational torsion subgroup of the variety increases as the dimension of the variety increases. In [8], fixing an integer $g\geq 2$, it was proved that any integer in the interval $[g^{2}+2g+1,g^{2}+3g+1]$ is realized as the order of a ${\mathbb{Q}}$-rational torsion point on the Jacobian of a hyperelliptic curve of genus $g$. Other families of hyperelliptic Jacobian varieties of dimension $g$ with ${\mathbb{Q}}$-rational torsion points of quadratic order in $g$ were given in [12, 15, 18]. In [7], it was conjectured that there exists a constant $\kappa$, independent of $g$, such that for every $m\leq\kappa g$, there exists an infinite family of hyperelliptic curves of genus $g$ over ${\mathbb{Q}}$ with a torsion point of order $m$ on its Jacobian. Leprévost proved this conjecture with $\kappa=3$ in [16].
@@
In this work, we extend the interval given by Leprévost to prove Flynn’s conjecture to include new integers that are realized as torsion orders for hyperelliptic Jacobian varieties. The divisor at infinity has finite order in the Jacobian of a hyperelliptic curve if and only if a certain continued fraction expansion of a power series is periodic. We show that for every integer $N$ in the interval $[3g,4g+1]$, $g\geq 3$, satisfying certain partition conditions, there exist infinite families of hyperelliptic curves of genus $g$ whose Jacobian varieties have a ${\mathbb{Q}}$-rational torsion point of order $N$. These partition conditions appear due to the fact that the order of the divisor at infinity in the Jacobian is the sum of the degrees of some partial quotients in the corresponding continued fraction expansion. Some of the integers satisfying these partition conditions are $14g+$ 1 when $g$ is odd; and $14g-$ 1 when $g$ is even. Consequently, we produce infinite families of hyperelliptic Jacobian varieties with torsion order $N=13$ when $g=3$, torsion $15$ when $g=4$, and torsion $17,18,21$ when $g=5$. These are the first examples of infinitely many varieties corresponding to the aforementioned pairs $(g,N)$. For a survey of all torsion points of a certain order that have been found for Jacobian varieties of hyperelliptic curves of genus $2,3$, and $4$, the reader may consult [23]. In addition, we show that infinitely many of the Jacobian varieties that we construct are absolutely simple.
@@ Acknowledgment
All the calculations in this work were performed using Magma, [3], and Mathematica, [25]. M. Sadek is supported by The Scientific and Technological Research Council of Turkey, TÜBİTAK, research grant ARDEB 1001/122F312.
@@ 2. Torsion divisors at infinity
Throughout this section $k$ is a field with $\operatorname{char}k\neq 2$. We let $\overline{k}$ be an algebraic closure of $k$.
2.1. Continued fractions in $K[x]$
Let $K=k(x)$ be function field over the field $k$. Define a function
$\varphi:K-k[x]\rightarrow K,\qquad\varphi(\alpha)=1/(\alpha-[\alpha])$
where $[\alpha]$ is the polynomial part of $\alpha$. For positive integers $r$ let $\varphi_{r}$ be the $r$-fold composition of $\varphi$ with itself and let $\varphi_{0}$ be the identity map, namely,
@@ 2.1. Continued fractions in $K[x]$
One can write the latter continued fraction in short as $\alpha=\left[a_{0},a_{1},a_{2},\ldots\right]$.
@@
An even length sequence $(a_{1},\cdots,a_{2r})$ is said to be skew symmetric of skew value $\gamma\in k-\{0\}$ if $a_{2(r-i)}=c_{i}a_{2i+1}$, $0\leq i<r$, where $c_{i}=\gamma$ if $i$ is even; and $c_{i}=1/\gamma$ otherwise.
@@ 2.2. The divisor at infinity
Let $C$ be a hyperelliptic curve of genus $g\geq 2$ over $k$ given by the equation $y^{2}=f(x)$ where $f(x)\in k[x]$ is of even degree $22g+$ 2, that has no repeated root in $\overline{k}$ and the leading coefficient of $f(x)$ is a nonzero square in $k$. Let $\infty_{+}$ and $\infty_{-}$ be the two $k$-rational points at infinity on $C$.
There exists a natural embedding of the curve $C$ into its Jacobian $J$ that maps a point $P$ to the divisor class $[P-D]$, where $D$ is a fixed divisor of degree $1$. This map restricts to $C(k)\xhookrightarrow{}J(k)$ if $D$ is a $k$-rational divisor. In particular, if there exists a point $P_{0}\in C(k)$, one can choose $D$ to be the rational divisor $P_{0}$.
The divisor at infinity on $C$ is the divisor $D_{\infty}=\infty_{+}-\infty_{-}$. It defines a point in $J(k)$. We say that $D_{\infty}$ is torsion of order $N$ if its class in $J$ has order $N$. The following result is due to Adams and Razar, see [1, Theorem 5.1].
@@
Let $C$ be a hyperelliptic curve described as above. The divisor $D_{\infty}$ is torsion if and only if the continued fraction of $y$ is periodic.
@@
    $y=[a_{0};\overline{a_{1},a_{2},\cdots,a_{m-1},2\gamma a_{0},a_{m-1},\cdots,a_{2},a_{1},2a_{0}}]$, where $(a_{1},a_{2},\cdots,a_{m-1})$ is skew symmetric of skew value $\gamma^{-1}$.
@@
The properties in Proposition 2.2 can be found in [1] and [24, Theorem 4.2].
@@
In Proposition 2.2, we call (the minimal) $m$ the quasi-period length. The (minimal) period length $n$ is equal to the quasi-period length $m$ exactly when $\gamma=1$. Further, if $m\neq n$, then $n=2m$ and $m$ must be odd. In the latter case, $m$ is called a strict quasi-period length.
@@ 3. Construction of curves
Fixing an integer $m\geq 0$, we set ${\mathbb{Q}}(\textbf{w}):={\mathbb{Q}}(w_{1},\cdots,w_{m})$, where $w_{1},\cdots,w_{m}$ are algebraically independent. In this section, we construct hyperelliptic curves $C$ of genus $g$ over ${\mathbb{Q}}(\textbf{w})$ described by the equation
$y^{2}=h_{2g+2}(\textbf{w})^{2}x^{2g+2}+h_{2g+1}(\textbf{w})x^{2g+1}+\cdots+h_{0}(\textbf{w}),\qquad h_{i}(\textbf{w})\in{\mathbb{Q}}(\textbf{w}),\quad i=0,\cdots,2g+2.$
The curve $C$ admits two ${\mathbb{Q}}(\textbf{w})$-rational points at infinity denoted by $\infty^{+}$ and $\infty^{-}$.
We investigate the case when the divisor $D_{\infty}=\infty^{+}-\infty^{-}$ in the Jacobian of $C$ is torsion with the quasi-period length $m$ is the same as period length $n$ and both are equal to $6$. This means that the continued fraction expansion of $y$ is given by
@@
In order to construct our desired curves, a solution of a certain polynomial Diophantine equation must be found.
@@
Let $r(x),q(x),a_{1}(x),a_{2}(x),a_{3}(x)\in{\mathbb{Q}}[x]$. The polynomial Diophantine equation
$-2q(x)-2a_{2}(x)+a_{1}(x)r(x)+a_{3}(x)r(x)-2q(x)a_{1}(x)a_{2}(x)+a_{1}(x)a_{2}(x)a_{3}(x)r(x)-a_{2}(x)^{2}a_{3}(x)\equiv 0$
has the following solution
$q(x)$ $=$ $u(x)r(x)/2,$
$a_{2}(x)$ $=$ $a_{1}(x)r(x),$
$a_{3}(x)$ $=$ $a_{1}(x)^{2}u(x)r(x)+a_{1}(x)+u(x),$
where $u(x)\in{\mathbb{Q}}[x]$.
@@
Fix an integer $g\geq 3$. Let $\alpha,\beta\geq 1,\gamma\geq 0$ be integers such that $12\alpha+2\beta+\gamma=g+$ 1. Let $a_{1}(x),r(x),u(x)\in{\mathbb{Q}}[x]$ be of degrees $\alpha,\beta,\gamma$, respectively. If the affine equation
$y^{2}=r(x)^{2}\left(u(x)(a_{1}(x)^{2}r(x)+1)+a_{1}(x)\right)^{2}+4\left(u(x)a_{1}(x)r^{2}(x)+r(x)\right)$
@@
$r(x)\left(u(x)(a_{1}(x)^{2}r(x)+1)+a_{1}(x)\right)^{2}+4\left(u(x)a_{1}(x)r(x)+1\right)$
is either the full symmetric group $\operatorname{\textbf{S}}_{2g+1}$ or the alternating group $\operatorname{\textbf{A}}_{2g+1}$, then the Jacobian $J_{C}$ of $C$ satisfies that $\operatorname{End}(J_{C})={\mathbb{Z}}$, in particular, $J_{C}$ is an absolutely simple abelian variety
Proof: In view of (1), a hyperelliptic curve $C$ for which the divisor at infinity is torsion with $m=n=6$ is described by an equation of the form
@@
In order for the latter equation to be a polynomial equation, the latter fraction must be a polynomial $h\in{\mathbb{Q}}[x]$. This yields that
@@
By comparing the degrees of the numerator and denominator of the latter fraction as polynomials in $a_{1},a_{2}$ and $a_{3}$, we may assume that $h=2qa_{2}+r$ with $\deg r<\deg q$ to obtain that
@@
We notice that if the degree of $q(x)$ is strictly less than the sum of the degrees of $a_{2}(x)$ and $a_{3}(x)$, then the numerator of the latter fraction is of strictly less degree than the denominator. Therefore, for $a_{0}$ to be a polynomial in ${\mathbb{Q}}[x]$, we may assume that $-2a_{2}-2q+a_{1}r+a_{3}r-2qa_{1}a_{2}+a_{1}a_{2}a_{3}r-a_{2}^{2}a_{3}\equiv 0\in{\mathbb{Q}}[x]$. According to Lemma 3.1, a solution to this polynomial Diophantine equation is provided giving rise to the affine equation $y^{2}=(a_{1}h+2q)^{2}/4+h$. After using the explicit description of the solution in Lemma 3.1, the curve $C$ is described by
$y^{2}=r(x)^{2}\left(u(x)(a_{1}(x)^{2}r(x)+1)+a_{1}(x)\right)^{2}+4\left(u(x)a_{1}(x)r^{2}(x)+r(x)\right).$
Now, when $\beta=1$, the rest of the statement will follow from [26, Theorem 1.3]. $\Box$
@@
We notice that in Theorem 3.2, if $a_{1}(x)=\sum_{i=0}^{\alpha}A_{i}x^{i}$, $r(x)=\sum_{i=1}^{\beta}R_{i}x^{i}$, and $u(x)=\sum_{i=0}^{\gamma}U_{i}x^{i}$, then the curve $C$ is defined by an equation of the form $y^{2}=f(x)$ where
@@
The discriminant of $C$ is a constant multiple of the discriminant $\Delta_{f}$ of the polynomial $f$, see for example [20] or [19], hence it is a polynomial in ${\mathbb{Q}}[A_{0},\cdots A_{\alpha},R_{0},\cdots R_{\beta},U_{0},\cdots,U_{\gamma}]$. If we consider the affine space $\mathbb{A}^{\alpha+\beta+\gamma+3}$ with coordinates $A_{0},\cdots A_{\alpha}$, $R_{0},\cdots R_{\beta}$, $U_{0},\cdots,U_{\gamma}$, then for a choice $(A_{0},\cdots A_{\alpha},R_{0},\cdots R_{\beta},U_{0},\cdots,U_{\gamma})\in\mathbb{A}^{\alpha+\beta+\gamma+3}$, the equation $y^{2}=f(x)$ defines a hyperelliptic curve of genus $g$ over ${\mathbb{Q}}$ if and only if $(A_{0},\cdots A_{\alpha},R_{0},\cdots R_{\beta},U_{0},\cdots,U_{\gamma})$ is chosen to lie in the complement of the hypersurface $\Delta_{f}=0$ in the affine space $\mathbb{A}^{\alpha+\beta+\gamma+3}$.
As a direct consequence of Theorem 3.2, we get the following result.
@@
Fix an integer $g\geq 3$. Let $N$ be an integer such that $13g\leq N\leq 4g+$ 1. If there exists integers $\alpha,\beta\geq 1,\gamma\geq 0$ such that $12\alpha+2\beta+\gamma=g+$ 1 and $N=g+1+6\alpha+3\beta+\gamma$, then there exists infinitely many hyperelliptic curves of genus $g$ over ${\mathbb{Q}}$ whose Jacobian posses a rational point of order $N$. In particular, if $g$ is odd, respectively even, there are infinitely many hyperelliptic curves of genus $g$ over ${\mathbb{Q}}$ whose Jacobian posses a rational point of order $14g+$ 1, respectively $14g-$ 1.
Although our trials to study hyperelliptic curves $C$ for which the corresponding continued fraction in Proposition 2.2 has quasi-period length $m\leq 5$ yielded torsion divisors at infinity with order $\operatorname{O}(kg)$, where $k<3$, we successfully obtained families of hyperelliptic curves whose jacobain contains torsion points whose order is not listed in literature. As an example, when $m=n=4$, one obtains the following equation
@@
Choosing $r,q$ to be ${\mathbb{Q}}$-rationals and $q^{2}\neq r$ when $g$ is odd; and $q$ to be ${\mathbb{Q}}$-rational whereas $r$ to be a linear polynomial if $g$ is even, the order of the torsion divisor at infinity becomes $25(g+1)/$ 2 if $g$ is odd; and $25g/2+$ 2 if $g$ is even.

### v2

@@ Abstract.
It was conjectured by Flynn that there exists a constant $\kappa$ such that, for any integer $g\geq 2$, any $m\leq\kappa g$, there exists a hyperelliptic curve of genus $g$ over $\mathbb{Q}$} with a rational $m$-torsion point on its Jacobian. Leprévost proved this conjecture with $\kappa=3$. In this work we prove that given an integer $N$ in the interval $[3g,4g+1]$, $g\geq 3$, satisfying certain partition conditions, there exist parametric families of hyperelliptic Jacobian varieties with a rational torsion point of order $N$. In particular, we establish the existence of such varieties for $N=4g+1$ when $g$ is odd and for N=4g1N=4g-1 when $g$ is even. A few explicit applications of this result produce the first known infinite examples of torsion $13$ when $g=3$, torsion $15$ when $g=4$, and torsion $17,18,21$ when $g=5$. In fact, we show that infinitely many of the latter abelian varieties are absolutely simple.
@@ 1. Introduction
A computational approach toward studying torsion subgroups of abelian varieties of dimension $d\geq 2$ is to find integers that are realized as orders of torsion points on these varieties. For example, a family of genus $2$ curves whose Jacobian varieties possess torsion points of order $\ell$ over $\mathbb{Q}$} was given in [8] when $\ell=11$, and in [14, 15, 17], when $\ell=13$, $15$, $17$, 1919 or $21$. In [10], several explicit families of curves of genus $2$ and $3$ whose Jacobians possess large rational torsion subgroups were displayed.
Researchers questioned the possibility of constructing families of abelian varieties in which the size of the rational torsion subgroup of the variety increases as the dimension of the variety increases. In [9], fixing an even integer $g\geq 2$, it was proved that any integer in the interval $g^{2}+2g+1,g^{2}+3g+1]$ is realized as the order of a $\mathbb{Q}$}-rational torsion point on the Jacobian of a hyperelliptic curve of genus $g$. Other families of hyperelliptic Jacobian varieties of dimension $g$ with $\mathbb{Q}$}-rational torsion points of quadratic order in $g$ were given in [13, 16, 19]. In [8], it was conjectured that there exists a constant $\kappa$, independent of $g$, such that for every $m\leq\kappa g$, there exists an infinite family of hyperelliptic curves of genus $g$ over $\mathbb{Q}$} with a torsion point of order $m$ on its Jacobian. Leprévost proved this conjecture with $\kappa=3$ in [18]. One may ask whether the integers between the linear bounds introduced by Leprévost and the quadratic bounds mentioned earlier are realizable as torsion orders of rational points of hyperelliptic Jacobian varieties. This manuscript attempts to address this gap in bounds and introduce new larger linear bounds.
@@
In this work, we extend the interval given by Leprévost to prove Flynn’s conjecture to include new integers that are realized as torsion orders for infinitely many hyperelliptic Jacobian varieties. The divisor at infinity has finite order in the Jacobian of a hyperelliptic curve if and only if a certain continued fraction expansion of a power series is periodic. We show that for every integer $N$ in the interval $[3g,4g+1]$, $g\geq 3$, satisfying certain partition conditions, there exist parametric families of hyperelliptic curves of genus $g$ whose Jacobian varieties have a $\mathbb{Q}$}-rational torsion point of order $N$. These partition conditions appear due to the fact that the order of the divisor at infinity in the Jacobian is the sum of the degrees of some partial quotients in the corresponding continued fraction expansion. Some of the integers satisfying these partition conditions are $4g+1$ when $g$ is odd; and 4g14g-1 when $g$ is even. Consequently, we produce infinite families of hyperelliptic Jacobian varieties with torsion order $N=13$ when $g=3$, torsion $15$ when $g=4$, and torsion $17,18,21$ when $g=5$. These are the first examples of infinitely many varieties corresponding to the aforementioned pairs $(g,N)$. For a survey of all torsion points of a certain order that have been found for Jacobian varieties of hyperelliptic curves of genus $2,3$, and $4$, the reader may consult [24]. In addition, we show that infinitely many of the Jacobian varieties that we construct are absolutely simple.
The idea that orders of torsion points on abelian varieties are linked to periodic continued fractions goes back to Abel and Chebychev. The work of Adams and Razar, [1], explained this link algebraically with much emphasis on torsion points of elliptic curves. However, continued fractions have not been used much to study torsion points on abelian varieties of higher dimension. We show that continued fractions over function fields can be employed to obtain new results on realizable torsion orders on Jacobian varieties of a fixed dimension.
@@ Acknowledgment
The authors would like to thank the anonymous referees for the thorough reading of the manuscript and for many useful suggestions and comments that improved the manuscript. All the calculations in this work were performed using Magma, [3], and Mathematica, [27]. This work builds upon the findings presented in the doctoral dissertation of Hamide Suluyer, [26]. M. Sadek is supported by The Scientific and Technological Research Council of Turkey, TÜBİTAK, research grant ARDEB 1001/122F312.
@@ 2. Torsion divisors at infinity
Throughout this section $k$ is a field with $\operatorname{char}k\neq 2$. We let $\overline{k}$ be an algebraic closure of $k$. The material in this section can be found in [1]. Let $C$ be a curve defined over a field $k$.
2.1. Continued fractions
Let $O\in C(k)$ be a non-singular $k$-rational point on $C$ and $K=k(C)$ be the corresponding function field. The valuation corresponding to $O$ will be denoted by $\operatorname{ord}_{O}$. Let $x\in K$ be a function with a simple pole at $O$. Then $x^{-1}$ is a uniformizing parameter at $O$. For any $\alpha\in K$, there is a unique $\beta\in k[x]$ such that $\operatorname{ord}_{O}(\alpha-\beta)>0$. We denote $\beta=[\alpha]$. We notice that the completion of $K$ at $O$ is the power series field $k((x^{-1}))$. If $\alpha$ is expanded as a Laurent series in $x^{-1}$, then $\alpha]$ is the non-positive part of the expansion, i.e., $\alpha]$ is a polynomial in $x$.
Define a function
$\begin{aligned} \varphi:K-k[x]\rightarrow K,\qquad\varphi(\alpha)=1/(\alpha-[\alpha])\end{aligned}.$
For positive integers $r$ let $\varphi_{r}$ be the $r$-fold composition of $\varphi$ with itself and let $\varphi_{0}$ be the identity map, namely,
@@ 2.1. Continued fractions
One can write the latter continued fraction in short as $\alpha=\left[a_{0};a_{1},a_{2},\ldots\right]$.
@@
Let $a_{i}$ be the $i$-th partial quotient of the continued fraction expansion of $y\in K$. The continued fraction of $y$ is said to be periodic if there is a positive integer $n$ such that $a_{h+n}=a_{h}$ for every $h\geq 1$. The smallest such $n$ is called the period of the continued fraction.
Definition 2.2.
A sequence $(a_{1},\cdots,a_{m-1})$ is said to be skew symmetric of skew value $\gamma\in k-\{0\}$ if $a_{(m-i)}=c_{i}a_{i}$, $1\leq i<(m+1)/2$, where $c_{i}=\gamma$ if $i$ is even; and $c_{i}=1/\gamma$ otherwise.
@@ 2.2. The divisor at infinity
Let $C$ be a hyperelliptic curve of genus $g\geq 2$ over $k$ given by the equation $y^{2}=f(x)$ where $f(x)\in k[x]$ is a separable polynomial of even degree $2g+2$, and the leading coefficient of $f(x)$ is a nonzero square in $k$. Let $\infty_{+}$ and $\infty_{-}$ be the two $k$-rational points at infinity on $C$.
There exists a natural embedding of the curve $C$ into its Jacobian $J$ that maps a point $P$ to the divisor class [PD][P-D], where $D$ is a fixed divisor of degree $1$. This map restricts to $C(k)\xhookrightarrow{}J(k)$ if $D$ is a $k$-rational divisor. In particular, if there exists a point $P_{0}\in C(k)$, one can choose $D$ to be the rational divisor $P_{0}$.
The divisor at infinity on $C$ is the divisor $D_{\infty}=\infty_{+}$-$\infty_{-}$. It defines a point in $J(k)$. We say that $D_{\infty}$ is torsion of order $N$ if its class in $J$ has order $N$. The following result can be found in [1, Theorem 5.1] when $g=1$, and [25, §4 and §5] when $g>1$.
@@
Let $C$ be a hyperelliptic curve described as above. The divisor $D_{\infty}$ is torsion if and only if the continued fraction of $\sqrt{f(x)}$ is periodic.
@@
    $\sqrt{f(x)}=[a_{0};\overline{a_{1},a_{2},\cdots,a_{m-1},2\gamma a_{0},a_{m-1},\cdots,a_{2},a_{1},2a_{0}}]$, where $(a_{1},a_{2},\cdots,a_{m-1})$ is skew symmetric of skew value $\gamma^{-1}$. If $m$ is even, then $\gamma=1$, hence the period of the continued fraction is $m$.
@@
In Proposition 2.3, we remark that the fact that $D_{\infty}$ is torsion gives rise to a unit $p(x)+q(x)\sqrt{f(x)}$ in $k(x,\sqrt{f(x)})$ where $p(x)/q(x)=[a_{0};a_{1},a_{2},\cdots,a_{m-1}]$. Moreover, the order of $D_{\infty}$ is the degree of $p(x)$. The interested reader may consult [25] for more details.
@@
In Proposition 2.3, we call (the minimal) $m$ the quasi-period length. Recalling that $a_{(m-i)}=a_{i}$, $1\leq i<(m+1)/2$, if $\gamma=1$, then the continued fraction expansion of $\sqrt{f(x)}$ is $a_{0};\overline{a_{1},a_{2},\cdots,a_{m-1},2a_{0},a_{1},\cdots,a_{m-1},2a_{0}}]$. One sees now that for $\gamma=1$, the (minimal) period length $n$ is equal to the quasi-period length $m$. Further, if $m\neq n$, then $n=2m$ and $m$ must be odd. In the latter case, $m$ is called a strict quasi-period length.
@@ 3. Construction of curves
Fixing an integer $r\geq 1$, we set ${\mathbb{Q}}(\textbf{w}):={\mathbb{Q}}(w_{1},\cdots,w_{r})$, where $w_{1},\cdots,w_{r}$ are indeterminates. In this section, we construct hyperelliptic curves $C$ of genus $g$ over ${\mathbb{Q}}(\textbf{w})$ described by the equation
$y^{2}=h_{2g+2}(\textbf{w})^{2}x^{2g+2}+h_{2g+1}(\textbf{w})x^{2g+1}+\cdots+h_{0}(\textbf{w}),\qquad h_{i}(\textbf{w})\in{\mathbb{Q}}(\textbf{w}),\quad i=0,\cdots,2g+2,$
where the discriminant of the latter equation is nonzero. The curve $C$ admits two ${\mathbb{Q}}(\textbf{w})$-rational points at infinity denoted by $\infty^{+}$ and $\infty^{-}$.
We investigate the case when the divisor $D_{\infty}=\infty^{+}$-$\infty^{-}$ in the Jacobian of $C$ is torsion with the quasi-period length $m$ and the period length $n$ are both equal to $6$. In view of Proposition 2.3, this implies that the continued fraction expansion of $y$ is given by
@@
The reason why we consider the case $m=n=6$ is that our attempts to construct hyperelliptic curves for which the corresponding quasi-period length $m\leq 5$ produced hyperelliptic curves whose Jacobians possess torsion divisors at infinity with order $O(kg)$, $k<3$, see [26, Chapter 3]. We recall that Leprévost, [18], established the existence of hyperelliptic curves of genus $g$ whose Jacobians possess rational torsion points of order $3g$.
The following lemma discusses the existence of polynomial solutions to a certain Diophantine equation. This equation will appear during the course of the proof of the main theorem of this section, Theorem 3.2, in which we construct our desired curves.
@@
Let $r,a_{1}\in\mathbb{Q}[x]$. Consider the following polynomial Diophantine equation
$2q-2a_{2}+a_{1}r+a_{3}r-2qa_{1}a_{2}+a_{1}a_{2}a_{3}r-a_{2}^{2}a_{3}=0.$
Then
$q$ $=\frac{ur}{2},$
$a_{2}$ $=a_{1}r,$
$a_{3}$ $=a_{1}^{2}ur+a_{1}+u,$
where $u\in{\mathbb{Q}}[x]$, provide a solution to this Diophantine equation.
@@
Fix an integer $g\geq 3$. Let $\alpha,\beta\geq 1,\gamma\geq 0$ be integers such that $2\alpha+2\beta+\gamma=g+1$. Let $a_{1},r,u\in{\mathbb{Q}}[x]$ be of degrees $\alpha,\beta,\gamma$, respectively. If the affine equation
$y^{2}=r^{2}\left(u(a_{1}^{2}r+1)+a_{1}\right)^{2}+4\left(ua_{1}r^{2}+r\right)$
@@
$r\left(u(a_{1}^{2}r+1)+a_{1}\right)^{2}+4\left(ua_{1}r+1\right)$
is either the full symmetric group $\operatorname{\textbf{S}}_{2g+1}$ or the alternating group $\operatorname{\textbf{A}}_{2g+1}$, then the Jacobian $J_{C}$ of $C$ satisfies that $\operatorname{End}(J_{C})={\mathbb{Z}}$, in particular, $J_{C}$ is an absolutely simple abelian variety.
Proof: In view of (1), a hyperelliptic curve $C$ for which the divisor at infinity is torsion with $m=n=6$ is described. Expanding the continued fraction (1) as a rational function yields that $C$ is described by an equation of the form
@@
In order for the latter equation to be a polynomial equation, the latter fraction must be a polynomial $h\in{\mathbb{Q}}[x]$, i.e., one has
$h=\frac{a_{2}(2+a_{2}a_{3})+2a_{0}(1+a_{2}a_{3}+a_{1}a_{2}(2+a_{2}a_{3}))}{(1+a_{1}a_{2})(a_{3}+a_{1}(2+a_{2}a_{3}))}\in{\mathbb{Q}}[x].$
This yields that
@@
We remark that if $\deg h\geq\deg a_{2}$, then the highest degree term of the numerator of the latter fraction is $a_{1}a_{2}a_{3}h$. In addition, the degree of the numerator will be larger than the degree of the denominator; which allows further division. Therefore, we assume that $h=2qa_{2}+r$ with $\deg r<\deg a_{2}$. We then obtain that
@@
We notice that if the degree of $q$ is strictly less than the sum of the degrees of $a_{2}$ and $a_{3}$, then the numerator of the latter fraction is of strictly less degree than the denominator. Therefore, for $a_{0}$ to be a polynomial in ${\mathbb{Q}}[x]$, we may assume that $2a_{2}-2q+a_{1}r+a_{3}r-2qa_{1}a_{2}+a_{1}a_{2}a_{3}r-a_{2}^{2}a_{3}=0\in{\mathbb{Q}}[x]$. According to Lemma 3.1, a solution to this polynomial Diophantine equation is provided giving rise to the affine equation $y^{2}=(a_{1}h+2q)^{2}/4+h$. After using the explicit description of the solution in Lemma 3.1, the curve $C$ is described by
$y^{2}=r^{2}\left(u(a_{1}^{2}r+1)+a_{1}\right)^{2}+4\left(ua_{1}r^{2}+r\right).$
Now, when $\beta=1$, the rest of the statement will follow from [28, Theorem 1.3]. $\Box$
@@
We notice that in Theorem 3.2, if $a_{1}=\sum_{i=0}^{\alpha}A_{i}x^{i}$, $r=\sum_{i=0}^{\beta}R_{i}x^{i}$, and $u=\sum_{i=0}^{\gamma}U_{i}x^{i}$, then the curve $C$ is defined by an equation of the form $y^{2}=f(x)$ where
@@
The discriminant of the equation $y^{2}=f(x)$ describing $C$ is a constant multiple of the discriminant $\Delta_{f}$ of the polynomial $f$, see for example [21] or [20], hence it is a polynomial in ${\mathbb{Q}}[A_{0},\cdots A_{\alpha},R_{0},\cdots R_{\beta},U_{0},\cdots,U_{\gamma}]$. If we consider the affine space $\mathbb{A}^{\alpha+\beta+\gamma+3}$ with coordinates $A_{0},\cdots A_{\alpha}$, $R_{0},\cdots R_{\beta}$, $U_{0},\cdots,U_{\gamma}$, then for a choice $(A_{0},\cdots A_{\alpha},R_{0},\cdots R_{\beta},U_{0},\cdots,U_{\gamma})\in\mathbb{A}^{\alpha+\beta+\gamma+3}$, the equation $y^{2}=f(x)$ defines a hyperelliptic curve of genus $g$ over $\mathbb{Q}$} if and only if $(A_{0},\cdots A_{\alpha},R_{0},\cdots R_{\beta},U_{0},\cdots,U_{\gamma})$ is chosen to lie in the complement of the hypersurface $\Delta_{f}=0$ in the affine space $\mathbb{A}^{\alpha+\beta+\gamma+3}$.
As a direct consequence of Theorem 3.2 and the paragraph above, we get the following result.
@@
Fix an integer $g\geq 3$. Let $N$ be an integer such that $3g\leq N\leq 4g+1$. If there exists integers $\alpha,\beta\geq 1,\gamma\geq 0$ such that $2\alpha+2\beta+\gamma=g+1$ and $N=g+1+6\alpha+3\beta+\gamma$, then there exists an $(\alpha+\beta+\gamma+3)$-parameter family of hyperelliptic curves of genus $g$ over $\mathbb{Q}$} whose Jacobians possess a rational point of order $N$. In particular, if $g$ is odd, respectively even, there is an $(\alpha+\beta+\gamma+3)$-parameter family of hyperelliptic curves of genus $g$ over $\mathbb{Q}$} whose Jacobians possess a rational point of order $4g+1$, respectively 4g14g-1.
Although our trials to study hyperelliptic curves $C$ for which the corresponding continued fraction in Proposition 2.3 has quasi-period length $m\leq 5$ yielded torsion divisors at infinity with order $\operatorname{O}(kg)$, where $k<3$, we successfully obtained families of hyperelliptic curves whose jacobain contains torsion points whose order is not listed in literature. As an example, when $m=n=4$, one has $y=[a_{0};\overline{a_{1},a_{2},a_{1},2a_{0}}]$. In other words, one sees that
$y^{2}=a_{0}^{2}+\frac{2a_{0}a_{1}a_{2}+2a_{0}+a_{2}}{2a_{1}+a_{1}^{2}a_{2}}.$
Now one may assume the existence of $h(x)\in k[x]$, $\deg h\leq g$, such that $h(2a_{1}+a_{1}^{2}a_{2})=2a_{0}a_{1}a_{2}+2a_{0}+a_{2}$. Therefore, one obtains
$a_{0}=\frac{h(2a_{1}+a_{1}^{2}a_{2})-a_{2}}{2(a_{2}a_{1}+1)}=\frac{ha_{1}}{2}+\frac{ha_{1}-a_{2}}{2(a_{1}a_{2}+1)}.$
Following the same ideas of the proof of Theorem 3.2, one may assume that $h=qa_{2}+r$, where $0\leq\deg q\leq g-1$ and $0\leq\deg r<\deg a_{2}$. Setting $a_{2}=ra_{1}-q$, one gets the equation
@@
Choosing $r,q$ to be $\mathbb{Q}$}-rationals and $q^{2}\neq r$ when $g$ is odd; and $q$ to be $\mathbb{Q}$}-rational whereas $r$ to be a linear polynomial if $g$ is even, the order of the torsion divisor at infinity becomes $5(g+1)/2$ if $g$ is odd; and $5g/2+2$ if $g$ is even, see Proposition 2.3.